# SAMG: State-Action-Aware Offline-to-Online Reinforcement Learning with Offline Model Guidance

### v1

@@ C.3. Datasets
C.4. SAMG performance
As illustrated in Section Empirical results, SAMG performs best when integrated with AWAC compared to other algorithms.
@@ C.4. SAMG performance
C.5. Cumulative regret
The cumulative regrets of the Antmaze environment of four vanilla algorithms and SAMG are shown in the Table 2.
@@ C.5. Cumulative regret
Table 2: Cumulative regret of online fine-tuning algorithms. The cumulative regret of standard base algorithms (including CQL, IQL, AWAC and Cal_QL, denoted as "Vanilla") compared to SAMG integrated algorithms (referred to as "Ours").
Dataset CQL AWAC IQL Cal-QL
@@
C.6. Unsatisfactory performance on particular environment
@@
C.7. Method to get the data coverage rate of offline dataset
To get the data coverage of a specific dataset, we aggregate all levels of datasets of a given environment, i.e., expert, medium-expert, medium-replay, medium, random level of datasets of environments HalfCheetah, Hopper, Walker2d. Thinking that the state and action are high-dimensional, we first perform dimensionality reduction. We uniformly and randomly select part of the data due to its huge scale and then perform t-SNE [Van der Maaten and Hinton, 2008] separately on the actions and states of this subset for dimensionality reduction. Given that it is hard to model the distribution of the continuous dimensional-reduced data, We then conduct hierarchical clustering [Müllner, 2011] to calculate and analyze the distribution of the data. We compute the clustering results of each environment and calculate the coverage rate based on the clustering results. To be specific, we select 10 percent of all data each time to cluster and repeat this process for 10 random seeds. For each clustering result, we calculate the data coverage rate of each level of offline dataset by counting the proportion of clustering center points. We consider one level of offline dataset to possess a clustering center if there exists more than 50 samples labeled with this clustering center.
Appendix D D. Compute resources

### v2

@@ C.3 Datasets
  5. 5
    Adroit: In the Adroit environment, the agent controls a robotic hand to finish various manipulation tasks, including pen balancing, door opening and object relocation. The agent only receives a positive reward when the task is successfully completed, otherwise, the reward is zero, which makes the tasks more challenging. We focus on three specific tasks: the pen agent must manipulate a pen to keep it balanced in some orientation; the door agent must grasp and open a door handle; the relocate agent must pick up an object and move it to a target location. We adopt the binary setting of Cal_QL [Nakamoto et al., 2023], which combines the cloned-level and human-level dataset. Therefore, there are tree different tasks: pen-binary, door-binary, relocate-binary.
C.4 SAMG performance
As illustrated in Section 5.1, SAMG performs best when integrated with AWAC compared to other algorithms.
@@ C.4 SAMG performance
C.5 Cumulative regret
The cumulative regrets of the Antmaze environment of four vanilla algorithms and SAMG are shown in Table 3.
@@ C.5 Cumulative regret
Table 3: Cumulative regret of online fine-tuning algorithms. The cumulative regret of standard base algorithms (including CQL, IQL, AWAC and Cal_QL, denoted as “Vanilla") compared to SAMG integrated algorithms (referred to as “Ours"). The result is the average normalized score of 5 random seeds $\pm$ (standard deviation). All algorithms are conducted for 500k iterations.
Dataset CQL AWAC IQL Cal_QL
@@
C.6 Unsatisfactory performance on particular environment
@@
C.7 Further comparisons with hybrid RL algorithms
We further compare the performance of Cal_QL with and without SAMG integrated. The results are illustrated in Table.
C.8 Method to get the data coverage rate of offline dataset
To get the data coverage of a specific dataset, we aggregate all levels of datasets of a given environment, i.e., expert, medium-expert, medium-replay, medium, random level of datasets of environments HalfCheetah, Hopper, Walker2d. Thinking that the state and action are high-dimensional, we first perform dimensionality reduction. We uniformly and randomly select part of the data due to its huge scale and then perform t-SNE [Van der Maaten and Hinton, 2008] separately on the actions and states of this subset for dimensionality reduction. Given that it is hard to model the distribution of the continuous dimensional-reduced data, We then conduct hierarchical clustering [Müllner, 2011] to calculate and analyze the distribution of the data. We compute the clustering results of each environment and calculate the coverage rate based on the clustering results. To be specific, we select 10 percent of all data each time to cluster and repeat this process for 10 random seeds. For each clustering result, we calculate the data coverage rate of each level of offline dataset by counting the proportion of clustering center points. We consider one level of offline dataset to possess a clustering center if there exist more than 50 samples labeled with this clustering center.
Appendix D Compute resources